# Quantum phase transition of (1+1)-dimensional O(3) nonlinear sigma model at finite density with tensor renormalization group

### v1

@@ 1 Introduction
In this paper we investigate the quantum phase transition of the (1+1)$d$ O(3) NLSM at finite density using the TRG method, which has been shown to be free from the sign problem or the complex action problem for various models Shimizu:2014uva ; Shimizu:2014fsa ; Kawauchi:2016xng ; Kawauchi:2016dcg ; Yang:2015rra ; Shimizu:2017onf ; Takeda:2014vwa ; Kadoh:2018hqq ; Kadoh:2019ube ; Kuramashi:2019cgs ; Akiyama:2021xxr ; Akiyama:2021glo ; Nakayama:2021iyp ; Luo:2022eje ; Akiyama:2020soe ; Akiyama:2020ntf ; Akiyama:2023hvt . We performed a detailed study of the phase tranasition at a fixed value of $\beta=1.4$, which is one of the inverse coupling employed in Ref. Luo:2023ont . The previous study with a collective Monte Carlo algorithm obtained the correlation length $\xi_{0}=6.90(1)$ at zero density Wolff:1989hv . The corresponding mass gap is $m=0.1449(2)$ so that the cutoff effects is expected to be small at $\beta=1.4$. We determine the transition point $\mu_{\rm c}$ and the critical exponent $\nu$ from the $\mu$ dependence of number density in the thermodynamic limit. The dynamical critical exponent $z$ is also determined from the $\mu$ dependence of the temporal correlation length $\xi_{t}$ extracted from the ratio of the first and the second eigen values in the transfer matrix.
@@ 2 Formulation and numerical algorithm
$T_{x_{n},x^{\prime}_{n},y_{n},y^{\prime}_{n}}$ $=\frac{\pi}{8}\sqrt{\sigma_{x_{n}}(1)\sigma_{x^{\prime}_{n}}(1)\sigma_{y_{n}}(2)\sigma_{y^{\prime}_{n}}(2)}\sum_{a_{n},b_{n}}w_{a_{n}}w_{b_{n}}$
@@ 3 Numerical results
The critical exponents $\nu$ and $z$ were previously studied using the dual lattice simulation Bruckmann:2016txt . Although the values of $\nu$ and $z$ were not directly determined, the authors observed that the assumption of $\nu=0.5$ and $z=2$ reasonably explains the results for the scaling properties of the spin stiffness. Our results of $\nu=0.512(15)$ and $z=1.96(1)$ confirm their observation.
@@ 4 Summary and outlook
We have investigated the properties of the quantum phase transition of the (1+1)$d$ O(3) NLSM at finite density with the TRG method. The critical chemical potential $\mu_{\rm c}=0.14512(11)$ at $\beta=1.4$ in the limit of $D_{\rm cut}\rightarrow\infty$ is consistent with the mass gap $m=1/\xi_{0}=1/6.90(1)=0.1449(2)$ at $\mu=0$ obtained in the previous work with a collective Monte Carlo algorithm Wolff:1989hv . We have also determined the critical exponent $\nu$ and the dynamical one $z$ for the transition. This is the first successful calculation of the dynamical critical exponent with the TRG method. We are now ready to investigate the properties of the quantum phase transitions of other finite density systems.
@@ Acknowledgements.
Numerical calculation for the present work was carried out with the supercomputers Cygnus and Pegasus under the Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba. This work is supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) (Nos. 24H00214, 24H00940).

### v2

@@ 1 Introduction
In this paper we investigate the quantum phase transition of the (1+1)$d$ O(3) NLSM at finite density using the TRG method, which has been shown to be free from the sign problem or the complex action problem for various models Shimizu:2014uva ; Shimizu:2014fsa ; Kawauchi:2016xng ; Kawauchi:2016dcg ; Yang:2015rra ; Shimizu:2017onf ; Takeda:2014vwa ; Kadoh:2018hqq ; Kadoh:2019ube ; Kuramashi:2019cgs ; Akiyama:2021xxr ; Akiyama:2021glo ; Nakayama:2021iyp ; Luo:2022eje ; Akiyama:2020soe ; Akiyama:2020ntf ; Akiyama:2023hvt . We performed a detailed study of the phase tranasition at a fixed value of $\beta=1.4$, which is one of the inverse coupling employed in Ref. Luo:2023ont . The previous study with a collective Monte Carlo algorithm obtained the correlation length $\xi_{0}=6.90(1)$ at zero density Wolff:1989hv . The corresponding mass gap is $m=0.1449(2)$ so that the cutoff effects is expected to be small at $\beta=1.4$. We determine the transition point $\mu_{\rm c}$ and the critical exponent $\nu$ from the $\mu$ dependence of number density in the thermodynamic limit. The dynamical critical exponent $z$ is also determined from the $\mu$ dependence of the temporal correlation length $\xi_{t}$ extracted from the ratio of the first and the second eigenvalues in the transfer matrix. Note that $\nu=0.5$ and $z=2$ are expected from the equivalence between the (1+1)$d$ O(3) NLSM at finite density and the integer-spin Heisenberg chain with a magnetic field, whose transition when changing the magnetic field belongs to the Pokrovsky-Talapov universality class Dzhaparidze_1978 ; Pokrovsky_1979 ; Schulz_1980 ; Schulz_1986 ; Affleck_1990 .
@@ 2 Formulation and numerical algorithm
$T_{x_{n},x^{\prime}_{n},y_{n},y^{\prime}_{n}}$ $=\frac{\pi}{8}\sqrt{\sigma_{x_{n}}(1)\sigma_{x^{\prime}_{n}}(1)\sigma_{y_{n}}(2)\sigma_{y^{\prime}_{n}}(2)}\sum_{a_{n},b_{n}}w_{a_{n}}w_{b_{n}}\sin(\theta_{a_{n}})$
@@ 3 Numerical results
The critical exponents $\nu$ and $z$ were previously studied using the dual lattice simulation Bruckmann:2016txt . Although the values of $\nu$ and $z$ were not directly determined, the authors observed that the assumption of $\nu=0.5$ and $z=2$ reasonably explains the results for the scaling properties of the spin stiffness. On the other hand, our results of $\nu=0.512(15)$ and $z=1.96(6)$ directly show the consistency with the theoretical expectation of $\nu=0.5$ and $z=2$.
@@ 4 Summary and outlook
We have investigated the properties of the quantum phase transition of the (1+1)$d$ O(3) NLSM at finite density with the TRG method. The critical chemical potential $\mu_{\rm c}=0.14512(11)$ at $\beta=1.4$ in the limit of $D_{\rm cut}\rightarrow\infty$ is consistent with the mass gap $m=1/\xi_{0}=1/6.90(1)=0.1449(2)$ at $\mu=0$ obtained in the previous work with a collective Monte Carlo algorithm Wolff:1989hv . We have also determined the critical exponent $\nu$ and the dynamical one $z$ for the transition. Our results of $\nu=0.512(15)$ and $z=1.96(6)$ are consistent with the theoretical expectation of $\nu=0.5$ and $z=2$. This is the first successful calculation of the dynamical critical exponent with the TRG method. We are now ready to investigate the properties of the quantum phase transitions of other finite density systems.
@@ Acknowledgements.
Numerical calculation for the present work was carried out with the supercomputers Cygnus and Pegasus under the Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba. We also used the supercomputer Fugaku provided by RIKEN through the HPCI System Research Project (Project ID: hp220203, hp230247). This work is supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) (Nos. 24H00214, 24H00940).